Evaluate the function at each specified value of the independent variable and simplify. (If an answer is undefined, enter UNDEFINED.)

g(y) = 7 − 5y
(a)
g(0) =

(b)
g






7


5





=
(c)
g(s + 1) =

Answer by stanbon(75887)

g(y) = 7 − 5y
(a)
g(0) = 7
---------------------
(b)
g(75) = 7-5*75 = -368
--------------------------------
(c)
g(s + 1) = 7-5(s+1) = 7-5s-5 = 2-5s
=========================================
Cheers,
Stan H.
-------------------